Starting at the origin, a beam of light hits a mirror(in the form of line) at the point A(4,8) and reflected line passes through the point B(8,12),  Compute the slope of the mirror
 

Grade:11

1 Answers

Vidhi Gupta
6 years ago
Slope of first line = 12-8/8-4 = 1Slope of second line = 8-0/4-0 = 2Tan(x) = m-1/m+1 (where m is Slope of the third line)Tan(x) = 2-m/1+2mSolving both the equations, We get 3m^2-2m-3 = 0Therefore m = 1+(10)^1/2 / 3